The equilibrium constant for the reaction NO(g) + O2(g) <=> 2NO2(g) decreases from 1.5 x 105 at 430ºC to 23 at 1000ºC. From these data, calculate the value of ΔHº for the reaction. ΔHº = _________ kJ/mol.
Answer #1

Nog + O2(g) E220029 K, = 2.5x105, K2 = 23 TE 430 +273 T2 = 1000 +273 = 1273 2 103 by using vant half equation log Kg = AH [I-